A voltaic cell is based on the following half reactions at 25 ⁰C:

Ag + + e- à Ag                                    E⁰ = 0.8V        

H2O2 + 2H+ + 2e- à 2H2O                   E⁰ = 1.78V      

Predict whether E cell is larger or smaller than E⁰ cell for the following cases. Explain your answers.

(a) [Ag +] = 1.0 M, [H2O2] = 2.0 M, [H+] = 2.0 M

(b) [Ag +] = 2.0 M, [H2O2] = 1.0 M, [H+] = 1.0 x 10-7 M
Answer #1

H2O2 + 2H+ +2 Ag \rightarrow 2 Ag+ + 2H2O

cell = 1.78V - 0.8V = 0.98V

(a)

Ecell = E°cell - RT/nF × ln Q

Ecell = 0.98 - 0.0591/2 × log((1.0)2/((2.0)2×2.0) = 1.007V

Ecell is greater than E°cell . (Answer)

(b)

Ecell = 0.98V - 0.0591/2 × log((2.0)2/((1.0×10-7)2×1.0)

= 0.55V

Ecell is smaller than the E°cell .